# No Other Representation Component Is Needed: Diffusion Transformers Can Provide Representation Guidance by Themselves

### v3

@@ Abstract
Recent studies have demonstrated that learning a meaningful internal representation can both accelerate generative training and enhance the generation quality of diffusion transformers. However, existing approaches necessitate to either introduce an external and complex representation training framework or rely on a large-scale, pre-trained representation foundation model to provide representation guidance during the original generative training process. In this study, we posit that the unique discriminative process inherent to diffusion transformers enables them to offer such guidance without requiring external representation components. We therefore propose Self-Representation Alignment (SRA), a simple yet straightforward method that obtains representation guidance through a self-distillation manner. Specifically, SRA aligns the output latent representation of the diffusion transformer in the earlier layer with higher noise to that in the later layer with lower noise to progressively enhance the overall representation learning during only the generative training process. Experimental results indicate that applying SRA to DiTs and SiTs yields consistent performance improvements. Moreover, SRA not only significantly outperforms approaches relying on auxiliary, complex representation training frameworks but also achieves performance comparable to methods that are heavily dependent on powerful external representation priors. Refer to caption Figure 1: Left: Methods like MaskDiT [84] and SD-DiT [86] use an external representation task to guide diffusion transformer. Middle: Methods like REPA [81] leverage an external representation foundation model as guidance. Right (our approach): We do not use any external representation component but still obtain such guidance through proposed self-representation alignment technique.
@@ 1 Introduction
Recently, many works [84, 86, 81, 42] have explored leveraging representation components of vision transformers for diffusion transformer’s training and have shown that learning a high-quality internal representation can not only speed up the generative training progress but also improve the generation quality. These works either utilize the training paradigm in representation learning (e.g., MAE’s [29], IBOT’s [85]) shown in Figure 1a or leverage a large-scale pre-trained representation foundation model (e.g., DINOv2 [61], CLIP [65]) shown in Figure 1b to give representation guidance for the diffusion transformer during the original generative training. However, the former method requires a complex training framework design, and the latter method relies on a powerful prior, which is trained on a massive amount of data with hundreds of GPUs. Thereout, an intriguing yet underexplored problem has come to light: Can we obtain such representation guidance only during the generative training without external representation components?
@@
Our observations: Different from the representation model that takes a clean image as input and then outputs semantically-rich feature, diffusion model often takes a noise latent as input and obtains cleaner one step-by-step. In other words, the generative mechanism by which the diffusion model operates can be generally considered as a coarse to fine process. Inspired by this behavior, we hypothesise that the representations in it also follow such a trend. To testify this, we perform an empirical analysis with recent diffusion transformers [56, 62]. As shown in Figure 2a, we first find out that the latent features in the diffusion transformer are progressively refined, moving from coarse to fine, as block layers increase and noise level decreases. Next, akin to the results in prior studies [81, 77], we observe that the diffusion transformer already learns meaningful discriminative representations as shown in Figure 2b. Meanwhile, although the accuracy drops off after reaching a peak at about layer 20 because the model needs to shift away to focus on generating images with high-frequency details, the quality of the representations basically transfers from bad to good by increasing block layers and decreasing noise level. These indicate that the diffusion transformer gets roughly from a coarse-to-fine discriminative process when only generative training is performed.
@@
Our approach: To overcome the abovementioned hurdles, we present Self-Representation Alignment (SRA), a simple but effective technique built on recent diffusion transformer architectures [56, 62]. As shown in Figure 1c, SRA does not need any representation component; in essence, it aligns the output latent representation in earlier layer conditioned with higher noise to that in later layer conditioned with lower noise to achieve self-representation alignment. Meanwhile, in order to make the training process more stable, we obtain the target features from another model that shares the same architecture with the trainable model but updates weight by weighted moving average (EMA)***In the following part, we use term ‘student’ to represent the trainable model and ‘teacher’ to represent the EMA model in SRA for simplicity.. Furthermore, the student’s output latent feature is first passed through the projection layers to conduct a slight nonlinear transformation for better representation extraction and then aligned with the target feature output by the teacher. In a nutshell, our SRA can offer a flexible way to integrate representation guidance without external component needs and architect modification.
@@
  • We introduce SRA, a simple yet effective method that aligns the output latent representation of the diffusion transformers in the earlier layer with higher noise to that in the later layer with lower noise to achieve self-representation alignment.
@@ 2.1 Preliminary: Training Object of DiT and SiT
Different from the denoise-based model, the flow-based model aims to learn a velocity field ${\mathbf{v}}_{\zeta}({\mathbf{x}}_{t},t)$ that governs a probability flow ordinary differential equation (PF ODE). This PF ODE allows the model to sample data by flowing toward the data distribution. This forward process of PF ODE is defined as:
@@ 2.2 Self-Representation Alignment
Figure 3: Overall framework of SRA. SRA explicitly aligns the diffusion transformer’s representation in the earlier layer with higher noise to that in the later layer with lower noise to achieve self-representation alignment.
@@ 3.2 Component-Wise Analysis
Table 1: Component-wise analysis on ImageNet 256$\times$256 without classifier-free guidance (CFG). $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. $m\xrightarrow{}n$ denotes aligning features from $m^{th}$ layer of the student with that from $n^{th}$ layer of the teacher. $0\sim k_{max}$ denotes time Interval $k$ is a random value in $[0,k_{max})$. PH. denotes whether to use the projection head.
@@
$106\xrightarrow{}$ 10 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 34.85 40.58
$84\xrightarrow{}$ 8 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 30.00 47.78
$104\xrightarrow{}$ 10 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 30.65 46.69
@@
$62\xrightarrow{}$ 6 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 32.14 46.36
$82\xrightarrow{}$ 8 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.31 50.13
$83\xrightarrow{}$ 8 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$33\xrightarrow{}$ 3 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 37.08 41.54
$83\xrightarrow{}$ 8 0 smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 31.07 47.32
@@
$83\xrightarrow{}$ 8 $0\sim 0.1$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.38 49.32
$83\xrightarrow{}$ 8 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 $0\sim 0.3$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.15 50.01
$83\xrightarrow{}$ 8 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 $0\sim 0.2$ $\ell_{1}$ Gen’s mom. $\checkmark$ 29.06 50.08
$83\xrightarrow{}$ 8 $0\sim 0.2$ $\ell_{2}$ Gen’s mom. $\checkmark$ 29.54 49.25
$83\xrightarrow{}$ 8 $0\sim 0.2$ smooth-$\ell_{1}$ Student copy $\checkmark$ 35.71 42.18
$83\xrightarrow{}$ 8 $0\sim 0.2$ smooth-$\ell_{1}$ SSL’s mom. $\checkmark$ 33.17 44.96
$83\xrightarrow{}$ 8 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\times$ 34.23 41.07
$83\xrightarrow{}$ 8 $0\sim 0.2$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
@@
Time interval for alignment. We then study the time interval (meaning the same as $k$ in Section 2.2) used for alignment. Here, we study fixed and dynamic intervals. Notably, we find that using the teacher’s features input with lower noise than the student’s leads to performance enhancement, and an interval value of 0.1 or the mean of 0.1 is optimal. We hypothesize that this is because lower noise levels can offer better representation guidance, but an excessively large time interval can hinder the model’s learning process, causing it to focus only on optimizing alignment loss at the expense of neglecting the generative aspects. As dynamic interval shows slightly better performance, we apply time interval as $0\sim 0.2$ in our feature experiments§§§For DiTs, we set time interval as $\lfloor 0\sim 200\rfloor$ by default since DiTs adopt the linear variance schedule with $t$ where $t\in\{0,1,2,...,999\}\cap\mathbb{Z}$..
@@ 3.3 System-Level Comparison
In this section, a system-level comparative study is performed to assess recent denoise-based approaches against diffusion transformers with SRA.
@@
Finally, we provide a quantitative comparison between SiT-XL with SRA and other recent dif- fusion model methods using classifier-free guidance [34]. Our method already outperforms the original SiT-XL model with 1000 fewer epochs, and it is further improved with longer training. At 800 epochs, SiT-XL with SRA achieves FID of 1.58 and IS of 311.4. It is worth noting that this result is far superior to methods (e.g., MaskDiT [84]) that rely on auxiliary and complex representation training frameworks and is comparable with methods (e.g., REPA [81]) that depend heavily on powerful external representation priors. Moreover, because of the progressively higher-quality guidance teacher provide throughout training, we find that our method is much less likely to encounter saturation compared with REPA (we provide detailed results and analysis in Appendix F).
@@ 4 Related Work
EMA model as teacher. Unlike traditional knowledge distillation [27] that uses a stronger pre-trained model as the teacher, using EMA model [36] as the teacher can be seen as self-distillation because the weight of the teacher model is obtained by the weighted moving average of the student. This method often needs a feasible pretext task to succeed. For example, MoCo [30, 14] sets the EMA teacher’s output as a momentum queue and uses contrastive learning to guide the student model; DINO [8, 61] feeds two views of images to the EMA teacher, and the trainable student then forces the student’s output distribution to be close to that of the teacher. Our work also shares some similarities, where we set aligning the student model’s latent feature in the earlier layer and higher noise with that in the later layer and lower noise of the EMA teacher as our pretext task for training.
@@ 5 Conclusion and Discussion
In this study, we show that diffusion transformers can provide representation guidance by themselves to boost generation performance with our proposed SRA, which aligns the latent representation of the diffusion transformer in the earlier layer with higher noise to that in the later layer with lower noise to progressively enhance the overall representation learning without external components.
@@ B.1 Denoising Diffusion Probabilistic Models (DiT)
Denoise denoise-based models [32, 60] aim to model the target distribution ( ${\mathbf{x}}\sim p({\mathbf{x}})$ ) by learning a gradual denoising process that transforms a Gaussian distribution ( $\mathcal{N}(\mathbf{0},\mathbf{I})$ ) into $p({\mathbf{x}})$. Formally, diffusion models learn a reverse process ( $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ ) corresponding to a pre-defined forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ ), which incrementally adds Gaussian noise to the data starting from $p({\mathbf{x}})$ over a sequence of time steps ( $t\in{1,\dots,T}$ ), with $T>0$ fixed.
@@ Appendix C Hyperparameter and More Implementation Details
Alignment time interval $0\sim 0.2$ $0\sim 0.2$ $0\sim 0.2$ $\lfloor 0\sim 200\rfloor$ $\lfloor 0\sim 200\rfloor$ $\lfloor 0\sim 200\rfloor$
@@ Appendix F Convergence Speed Against REPA
As shown in Figure 8, although REPA converges quickly at the beginning of training due to its strong representation prior, the performance saturates after about 200 epochs. On the contrary, in SRA, since the teacher’s ability is constantly enhanced during training and thus better and better guidance is provided, continuous performance improvement can be achieved. From another perspective, the representation guidance signal in REPA is fixed in every timestep and epoch, while our supervision signal is dynamically changing. This provides a wider variety of learning cues to help the model continuously improve its performance.

### v4

@@ Abstract
Recent studies have demonstrated that learning a meaningful internal representation can both accelerate generative training and enhance the generation quality of diffusion transformers. However, existing approaches necessitate to either introduce an external and complex representation training framework or rely on a large-scale, pre-trained representation model to provide representation guidance during the original generative training process. In this study, we posit that the unique discriminative process inherent to diffusion transformers enables them to offer such guidance without requiring external representation components. We therefore propose Self-Representation Alignment (SRA), a simple yet effective method that obtains representation guidance through a self-distillation manner. Specifically, SRA aligns the latent representation of the diffusion transformer in the earlier layer conditioned on higher noise to that in the later layer conditioned on lower noise to progressively enhance the overall representation learning during only the generative training process. Experimental results indicate that applying SRA to DiTs and SiTs yields consistent performance improvements. Moreover, SRA not only largely outperforms approaches relying on auxiliary, complex representation training frameworks but also achieves performance comparable to methods that are heavily dependent on an external pre-trained representation model. Refer to caption Figure 1: Left: Methods like MaskDiT [84] and SD-DiT [86] use an external representation task to guide diffusion transformer. Middle: Methods like REPA [81] leverage an external representation foundation model as guidance. Right (our approach): We do not use any external representation component but still obtain such guidance through proposed self-representation alignment technique.
@@ 1 Introduction
Recently, many works [84, 86, 81, 42] have explored leveraging representation components of vision transformers for diffusion transformer’s training and have shown that learning a high-quality internal representation can not only speed up the generative training progress but also improve the generation quality. These works either utilize the training paradigm in representation learning (e.g., MAE’s [29], IBOT’s [85]) shown in Figure 1a or leverage a large-scale pre-trained representation foundation model (e.g., DINOv2 [61], CLIP [65]) shown in Figure 1b to give representation guidance for the diffusion transformer during the original generative training. However, the former method requires a complex training framework design, and the latter method relies on a powerful model, which is trained on a massive amount of data with hundreds of GPUs. Thereout, an intriguing yet underexplored problem has come to light: Can we obtain such representation guidance only during the generative training without external representation components?
@@
Our observations: Different from the representation model that takes a clean image as input and then outputs semantically-rich feature, diffusion model often takes a noise latent as input and obtains cleaner one step-by-step. In other words, the generative mechanism by which the diffusion model operates can be generally considered as a coarse to fine process. Inspired by this behavior, we hypothesise that the representations in it also follow such a trend. To testify this, we perform an empirical analysis with recent diffusion transformers [56, 62]. As shown in Figure 2a, we first find out that the latent features in the diffusion transformer are progressively refined, moving from coarse to fine, as block layers increase and noise level decreases. Next, akin to the results in previous studies [81, 77], we observe that the diffusion transformer already learns meaningful discriminative representations as shown in Figure 2b. Meanwhile, although the accuracy drops off after reaching a peak at about layer 20 because the model needs to shift away to focus on generating images with high-frequency details, the quality of the representations basically transfers from bad to good by increasing block layers and decreasing noise level. These indicate that the diffusion transformer gets roughly from a coarse-to-fine discriminative process when only generative training is performed.
@@
Our approach: To overcome the abovementioned hurdles, we present Self-Representation Alignment (SRA), a simple but effective technique built on recent diffusion transformer architectures [56, 62]. As shown in Figure 1c, SRA does not need any representation component; in essence, it aligns the output latent representation in earlier layer conditioned on higher noise to that in later layer conditioned on lower noise to achieve self-representation alignment. Meanwhile, in order to make the training process more stable, we obtain the target features from another model that shares the same architecture with the trainable model but updates weight by weighted moving average (EMA)***In the following part, we use term ‘student’ to represent the trainable model and ‘teacher’ to represent the EMA model in SRA for simplicity.. Furthermore, the student’s output latent feature is first passed through the projection layers to conduct a slight nonlinear transformation for better representation extraction and then aligned with the target feature output by the teacher. In a nutshell, our SRA can offer a flexible way to integrate representation guidance without external component needs and architect modification.
@@
  • We introduce SRA, a simple yet effective method that aligns the output latent representation of the diffusion transformers in the earlier layer conditioned on higher noise to that in the later layer conditioned on lower noise to achieve self-representation alignment.
@@ 2.1 Preliminary: Training Object of DiT and SiT
Different from the denoise-based model, the flow-based model aims to learn a velocity field ${\mathbf{v}}_{\zeta}({\mathbf{x}}_{t},t)$ that governs a probability flow ordinary differential equation (PF ODE). This PF ODE allows the model to sample data by flowing toward the data distribution. This forward process is defined as:
@@ 2.2 Self-Representation Alignment
Figure 3: Overall framework. SRA aligns the student’s latent representation in the earlier layer conditioned on higher noise (green branch) to that of the teacher in the later layer conditioned on lower noise (blue branch) to achieve self-representation alignment. We use a stop-gradient (sg) operator on the teacher to let gradients flow only through the student, and update the teacher’s parameters with an exponential moving average (ema) of the student’s parameters.
@@ 3.2 Component-Wise Analysis
Table 1: Component-wise analysis on ImageNet 256$\times$256 without classifier-free guidance (CFG). $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. $m\xrightarrow{}n$ denotes aligning features from $m^{th}$ layer of the student with that from $n^{th}$ layer of the teacher. $[0,k_{max})$ denotes time interval $k$ is chosen randomly from 0 to $k_{max}$. PH. denotes whether to use the projection head.
@@
$106\xrightarrow{}$ 10 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 34.85 40.58
$84\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 30.00 47.78
$104\xrightarrow{}$ 10 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 30.65 46.69
@@
$62\xrightarrow{}$ 6 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 32.14 46.36
$82\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.31 50.13
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$33\xrightarrow{}$ 3 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 37.08 41.54
$83\xrightarrow{}$ 8 0.0 smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 31.07 47.32
@@
$83\xrightarrow{}$ 8 0$[0,0.1)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.38 49.32
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 0$[0,0.3)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.15 50.01
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 0$[0,0.2)$ $\ell_{1}$ Gen’s mom. $\checkmark$ 29.06 50.08
$83\xrightarrow{}$ 8 0$[0,0.2)$ $\ell_{2}$ Gen’s mom. $\checkmark$ 29.54 49.25
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Student copy $\checkmark$ 35.71 42.18
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ SSL’s mom. $\checkmark$ 33.17 44.96
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\times$ 34.23 41.07
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
@@
Time interval for alignment. We then study the time interval (meaning the same as $k$ in Section 2.2) used for alignment. Here, we study fixed and dynamic intervals. Notably, we find that using the teacher’s features input with lower noise level than the student’s leads to performance enhancement, and an interval value of 0.1 or the mean of 0.1 is optimal. We hypothesize that this is because lower noise levels can offer better representation guidance, but an excessively large time interval can hinder the model’s learning process, causing it to focus only on optimizing alignment loss at the expense of neglecting the generative aspects. As dynamic interval shows slightly better performance, we apply time interval as $0\sim 0.2$ in our feature experiments§§§For DiTs, we set time interval as $\lfloor[0,200)\rfloor$ by default since DiTs adopt the linear variance schedule with $t$ where $t\in\{0,1,2,...,999\}\cap\mathbb{Z}$..
@@ 3.3 System-Level Comparison
In this section, a system-level comparative study is performed to assess recent diffusion-based approaches against diffusion transformers with SRA.
@@
Finally, we provide a quantitative comparison between SiT-XL with SRA and other recent diffusion model methods using classifier-free guidance [34]. Our method already outperforms the original SiT-XL model with 1000 fewer epochs, and it is further improved with longer training. At 800 epochs, SiT-XL with SRA achieves FID of 1.58 and IS of 311.4. It is worth noting that this result is far superior to methods (e.g., MaskDiT [84]) that rely on auxiliary and complex representation training frameworks and is comparable with methods (e.g., REPA [81]) that depend heavily on an external pre-trained representation model. Moreover, because of the progressively higher-quality guidance teacher provide throughout training, we find that our method is much less likely to encounter saturation compared with REPA (we provide detailed results and analysis in Appendix F).
@@ 4 Related Work
EMA model as teacher. Unlike traditional knowledge distillation [27] that uses a well pre-trained model as the teacher, using EMA model [36] as the teacher can be seen as self-distillation because the weight of the teacher model is obtained by the weighted moving average of the student. This method often needs a feasible pretext task to succeed. For example, MoCo [30, 14] sets the teacher’s output as a momentum queue and uses contrastive learning to guide the student model; DINO [8, 61] feeds two views of images to the teacher, and the trainable student then forces the student’s output distribution to be close to that of the teacher. Our work also shares some similarities, where we set aligning the student model’s latent feature in the earlier layer conditioned on higher noise with that in the later layer conditioned on lower noise of the teacher as our pretext task for training.
@@ 5 Conclusion and Discussion
In this study, we show that diffusion transformers can provide representation guidance by themselves to boost generation performance with our proposed SRA, which aligns their latent representation in the earlier layer conditioned on higher noise to that in the later layer conditioned on lower noise to progressively enhance the representation learning without external components.
@@ B.1 Denoising Diffusion Probabilistic Models (DiT)
Denoise-based models [32, 60] aim to model the target distribution ( ${\mathbf{x}}\sim p({\mathbf{x}})$ ) by learning a gradual denoising process that transforms a Gaussian distribution ( $\mathcal{N}(\mathbf{0},\mathbf{I})$ ) into $p({\mathbf{x}})$. Formally, diffusion models learn a reverse process ( $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ ) corresponding to a pre-defined forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ ), which incrementally adds Gaussian noise to the data starting from $p({\mathbf{x}})$ over a sequence of time steps ( $t\in{1,\dots,T}$ ), with $T>0$ fixed.
@@ Appendix C Hyperparameter and More Implementation Details
Alignment time interval $[0,0.2)$ $[0,0.2)$ $[0,0.2)$ $\lfloor[0,200)\rfloor$ $\lfloor[0,200)\rfloor$ $\lfloor[0,200)\rfloor$
@@ Appendix F Convergence Speed Against REPA
As shown in Figure 8, although REPA converges quickly at the beginning of training due to its large-scale, pre-trained representation model, the performance saturates after about 200 epochs. On the contrary, in SRA, since the teacher’s ability is constantly enhanced during training and thus better and better guidance is provided, continuous performance improvement can be achieved. From another perspective, the representation guidance signal in REPA is fixed in every timestep and epoch, while our supervision signal is dynamically changing. This provides a wider variety of learning cues to help the model continuously improve its performance.